OpenStudy (anonymous):

The 9 officers of the Student Council are going on a trip to an amusement park. Each student must pay an entrance fee plus $5 for meals. The total cost of the trip is $225. Solve the equation

OpenStudy (apoorvk):

let entrance fee per kid be x total cost per kid including meals = x+5 total cost for 9 kids = 9(x+5) total cost actually = 225 so, 225 = 9(x+5) --> now solve for 'x'.